OpenStudy (anonymous):

Someone please help. |3t+6|=9

OpenStudy (anonymous):

Do you know what those lines mean?

OpenStudy (anonymous):

no not really.

OpenStudy (the_fizicx99):

Subtract 6 from both sides then divide by 3. Then make the 9 negative, -9 and do the same process again.

OpenStudy (shamil98):

|3t+6|=9 you solve two equations 3t + 6 = 9 and 3t + 6 = -9

OpenStudy (anonymous):

well everything inside of it whether it is negative is going to result in a positive number.

OpenStudy (the_fizicx99):

=.= stole ma thunder bruh^^

OpenStudy (anonymous):

is the answer 1,-1

OpenStudy (the_fizicx99):

Nope

OpenStudy (the_fizicx99):

-9-6 = ?

OpenStudy (anonymous):

got it. 1, -5 ??

OpenStudy (the_fizicx99):

Yup